OpenStudy (anonymous):

A voltage source e(t), which has a value of 2V at t=0, rises exponentially to a final value of 6V with a time constant of 3sec. Determine the expression for e(t).

OpenStudy (anonymous):

e(0) = 2V e(3) = 6V = 3 * 2V = 3 * e(0) e(t) = t * e(0)

OpenStudy (anonymous):

I have possible answers a) e(t)=(6-4e^(-t/3))u(t) b) e(t)=6e^(-t/3)u(t) c)e(t)=(2+4e^(-t/3))u(t) d)e(t)(6-2e^(-3t))u(t)

OpenStudy (anonymous):

Thank you!